OpenStudy (king2001):

Identify the horizontal middle of a sample period for the function: f (t) = 4sec(3t − 6π). I got the answer x=4pi/3

OpenStudy (abdullahm):

Do you have answer choices for this question? There are infinite answers to this question.

OpenStudy (king2001):

the answer choices are x=4pi/3, x=5pi/2, x=pi, and x=3+pi

OpenStudy (abdullahm):

x = pi and x = 4pi/3 both work. http://prntscr.com/ayhmg3

OpenStudy (king2001):

yes but there is only 1 choice

OpenStudy (math&ing001):

y = f(x) = a cos ( k( x - b) ) + c Where : "a" is the amplitude "b" is the horizontal shift The horizontal shift here would be \(2\pi\) So I'm guessing what they mean by horizontal middle is half of the horizontal shift \(\pi\).